# Arecibo Wow! I: An Astrophysical Explanation for the Wow! Signal

### v1

@@ 3 SETI efforts from Arecibo
Following the end of HRMS, Arecibo continued to participate in SETI through various initiatives, including the Search for Extraterrestrial Radio Emissions from Nearby Developed Intelligent Populations (SERENDIP) project (Werthimer et al., 1988). SERENDIP was a piggyback search, which means that it used data collected by Arecibo during other observations to perform a parallel search for extraterrestrial signals. SERENDIP IV, which began in the late 1990s, used a multichannel receiver installed at Arecibo to scan the sky for signals over a wide range of frequencies (Chennamangalam et al., 2017). The data collected by SERENDIP IV were also used by the SETI@home project, a distributed computing effort that allowed millions of volunteers around the world to analyze SETI data on their home computers (Korpela et al., 2001). This collaboration between Arecibo and SETI@home significantly expanded the scope and reach of SETI research.
@@
Throughout its operational life, Arecibo was involved in numerous other SETI-related projects and observations, often serving as a testbed for new technologies and techniques in the search for extraterrestrial intelligence. The observatory’s unparalleled sensitivity and versatility made it a cornerstone of SETI research, and its contributions continue to influence the field even after its collapse in 2020. The Arecibo Observatory’s legacy in SETI is marked by its role in pioneering early searches, conducting high-profile experiments like the Arecibo Message, and supporting a wide range of projects that advanced our understanding of the cosmos and the potential for life beyond Earth. The observatory data and findings remain valuable resources for ongoing and future SETI efforts (Tarter, 2001). The Big Ear and the Arecibo Observatory were two contemporary telescopes that dedicated some of their time to SETI (see Table 2).
@@ 4 Arecibo REDS and Wow! Projects
In 2017 we started the Arecibo REDS project (Radio Emissions from Red Dwarf Stars). This project aims to study radio emissions from red dwarf stars using data from the Arecibo Observatory. Red dwarf stars, the most abundant type in the Milky Way, emit a wide range of radio frequencies because of their high activity levels (Yiu et al., 2024). We are studying these emissions, focusing on their frequency, polarization, and variability. Our observations are intended to help characterize their stellar flares and coronal mass ejections, which impact the space weather and habitability of their planets (Airapetian et al., 2020). Our research also explores the potential utility of these observations as tools for detecting exoplanets orbiting red dwarf stars (Vedantham et al., 2020). The results of Arecibo REDS will be presented elsewhere. After the collapse of the Arecibo Telescope in 2020, we plan to continue these observations with other radio telescopes.
@@
This study constitutes an extension of our Arecibo REDS project. It describes the initial findings of our Arecibo Wow! technosignature project. We observed from the Arecibo Observatory’s 305-meter telescope from 2017 to 2020 (Project A3123) and the 12-meter telescope in 2023 (Project A4006). Most of our targets were red dwarf stars with planets. Our observation protocol was similar to the one used by Route & Wolszczan (2012) to search for radio flares from ultracool dwarfs. Spectral observations were made in the L-band in full Stokes mode with up to 10-minute integration using the Arecibo Observatory’s Mock Spectrometer. Each of the seven receivers used had a bandwidth of 83 MHz at central frequencies of 1282, 1353, 1424, 1495, 1566, 1637, and 1708 MHz, for a total continuous bandwidth of 509 MHz. Similar observations were made in the S, C, and X bands. The integration was 100 ms and the channel width 10 kHz (unintentionally the same as Big Ear during the Wow! Signal signal detection), distributed in 8192 channels per receiver. We used both diode and target calibrators to calibrate our flux density measurements.
@@
Our observational strategy was neither tailored nor optimized for the detection of technosignatures. However, our sample included stars with known potentially habitable exoplanets, which were observed for long sessions spanning frequencies from 1 to 10 GHz. In this study, we focus only on our drift-scan mode observations made from February to May 2020 at topocentric frequencies between 1419.5 and 1421.0 MHz (see Table 5). This mode and frequency range are analogous to the protocol used in the detection of the Wow! Signal (see Table 3). Coincidentally, our 10 kHz channel width was the same. The seven target samples included Teegarden’s Star, Barnard’s Star, and HD 157881, which were part of our main study (see Table 4). We also conducted observations in this mode of the radio galaxies 3C 76.1 (B0300+162) and 3C 365 (B1756+134), which served as our calibration standards. The magnetar SGR 1935+2154 constituted an opportunistic target, due to its active state during the observation period and its proximal spatial alignment with our primary subjects (Bochenek et al., 2020). Our final target was a nearby random section of the sky that we denoted Void, for reference purposes.
Table 5: Observation dates pertinent to this study. The table enumerates only those targets observed in drift mode and within the L-band spectrum. In addition, these and other targets were observed in varying modes, including calibrators. The date (AST) and scan ID columns are provided to facilitate the identification of the Arecibo FITS files with these data sets. These observations are part of the Arecibo Observatory project A3123.
Date (AST) Date (MJD) RA (deg) DEC (deg) Target Scan ID Time (min)
20200210 58889.960405 43.1284 16.8811 Teegarden 52 1
20200225 58904.852072 42.7526 16.8806 Teegarden 18 4
58904.886898 42.0009 16.8797 Teegarden 33 10
58904.895660 45.1845 16.4455 B0300+162 36 5
58904.901667 42.6273 16.8805 Teegarden 39 5
58904.914306 44.3753 16.9992 Void 52 5
58904.920046 42.6290 16.8805 Teegarden 55 5
20200430 58969.297257 269.3267 4.6936 Barnard 18 1
58969.298576 269.3267 4.6936 Barnard 21 1
58969.299896 269.3267 4.6936 Barnard 24 1
58969.361435 293.6067 21.8969 SGR1935+2154 62 1
58969.362789 293.6067 21.8969 SGR1935+2154 65 1
20200506 58975.263692 269.5042 13.4759 B1756+134 6 1
58975.270880 261.3131 2.1117 HD157881 13 1
58975.279387 269.3267 4.6936 Barnard 24 1
58975.344491 293.6067 21.8969 SGR1935+2154 60 1
58975.349734 293.6067 21.9219 SGR1935+2154 67 1
58975.351389 293.6067 21.8969 SGR1935+2154 70 1
58975.352697 293.6067 21.9219 SGR1935+2154 73 1
58975.354005 293.6068 21.9469 SGR1935+2154 76 1
20200513 58982.307755 269.3268 4.5769 Barnard 37 1
58982.309086 269.3269 4.6061 Barnard 40 1
58982.310417 269.3270 4.6353 Barnard 43 1
58982.311748 269.3270 4.6644 Barnard 46 1
58982.313067 269.3271 4.6936 Barnard 49 1
58982.314398 269.3271 4.7228 Barnard 52 1
58982.315718 269.3272 4.7519 Barnard 55 1
58982.317037 269.3272 4.7811 Barnard 58 1
58982.322338 293.6068 21.7803 SGR1935+2154 61 1
58982.323646 293.6068 21.8094 SGR1935+2154 64 1
58982.324954 293.6069 21.8386 SGR1935+2154 67 1
58982.326262 293.6070 21.8678 SGR1935+2154 70 1
58982.327569 293.6070 21.8969 SGR1935+2154 73 1
58982.328877 293.6071 21.9261 SGR1935+2154 76 1
58982.330185 293.6071 21.9553 SGR1935+2154 79 1
58982.331493 293.6072 21.9844 SGR1935+2154 82 1
@@ 5 Data Analysis and Results
We devised a series of methodologies in Interactive Data Language (IDL) aimed at the detection of narrowband signals. The data were read and calibrated using the Arecibo Software Library using standard procedures. Among these methodologies, one closely emulated the analytical approach employed by the Big Ear Observatory to detect the Wow! Signal, specifically averaging data over 10-second intervals and utilizing the most recent 60 data points for background noise estimation. All of the methods yielded consistent results.
@@
The dynamic spectra were detrended and normalized in frequency and time to cancel out broadband signals and identify narrowband variations. The signals were converted to signal-to-noise ratio (SNR) to mimic the analysis used by the Big Ear to detect the Wow! Signal. Most of the observations were made in the LSRK frame but were converted to the topocentric frame to make it easier to identify terrestrial RFI. Our longest drift scan was 10 minutes long (2 degrees in right ascension), of Teegarden’s star (see Figure 2).
@@
Figure 2: Dynamic spectra of a 10-minute drift scan centered on Teegarden’s Star. The top subplot is the average flux density as a function of frequency showing the complete galactic HI profile. The right subplot shows a peak between 400 and 500 seconds attributable to AGC 74887. At least, four narrowband signals, with one reaching a seven-sigma peak, are observed (labeled A to D, from top to bottom). The signals are not associated with either Teegarden’s Star or AGC 74887. The observations were conducted in the LSRK frame, whereas the plot is rendered in the topocentric frame.
@@
We identified four main signals in the vicinity of Teegarden’s Star corresponding to small angular-size cold hydrogen clouds (see Figure 2). Similar clouds appeared in the observations of SGR 1935+2154 and in the Void regions (see Table 4). The intensity and location of the cloud signals were consistent between multiple independent scans during the same observation session of Teegarden’s Star. There was some variability for SGR 1935+2154, which included three observation sessions separated by a week.
@@
We focus our analysis in the clouds near Teegarden’s Star since it was the longest drift scan. Their observed properties are consistent with cold hydrogen clouds, assuming that they are optically thin (see Table 6). Their hydrogen column density is their partial contribution to the HI background (see Figure 3). All the results are available in this url.
The most noticeable characteristic of the signal depicted in Figure 2, is the close resemblance of the dynamic spectra of the observed region with the Wow! Signal (see Figure 1). At low spatial resolution the HI clouds in the line-of-sight produce the characteristic profile of neutral hydrogen emission (see the integrated spectrum in the upper panel). At the higher resolution of Arecibo (and Big Ear), what appears to be a single cloud reveals itself as several spatially resolved sources having different emission frequencies (due to different velocities).
Although the angular sizes of the observed clouds in the Arecibo signals are comparable to the azimuth beam width of the Big Ear telescope, namely 8 arcminutes (see Table 3), their intensities are below the detection limit of the Big Ear radio telescope. They also share the same narrowband frequency bandwidth of about 10 kHz or less. The only difference between the signals observed in Arecibo and the Wow! Signal is their brightness. It is precisely the similarity between these spectra that suggests a mechanism for the origin of the mysterious signal.
6 A mechanism for the Wow! Signal
We postulate that when observed by the Big Ear telescope in 1977 one of the clouds in the beam of the telescope was experiencing a transient increase in brightness due to a natural astrophysical process. Numerous well-known mechanisms could account for this phenomenon; however, the majority of such processes are also likely to increase the thermal energy or kinematic activity of hydrogen within the cloud, thereby broadening the hydrogen spectral line beyond the 10 kHz range. The only mechanism able to produce such a narrow-band signal is a maser.
Astronomical hydrogen masers are rare. The first one was discovered in 1989 around the star MWC349 (Martin-Pintado et al., 1989). Its emission exhibited characteristics similar to those of a molecular maser source. It was highly luminous and fluctuated over time, because of its sensitivity to alterations in the intricate excitation processes.
Hydrogen masers operating at the HI line frequency have yet to be detected in astronomical observations, though they have been produced in laboratory settings (Major & Major, 2007). In fact, human-made hydrogen masers are used for timekeeping in astronomical observatories (Gray, 2012). This experimental evidence hints that there are no fundamental limitations for which hydrogen in the neutral medium could be pumped and stimulated to produce maser-like emission.
The cold neutral medium of the Interstellar Medium (ISM) is predominantly composed of atomic hydrogen (HI), which exists primarily as clouds dispersed throughout the galaxy. These clouds are characterized by temperatures ranging from 50 to 100 K and densities between 10 and 100 atoms per cubic centimeter. The equilibrium brightness of an HI cloud, governed primarily by thermal emission processes, is relatively low, depending on the temperature and density of the cloud. However, if a maser were active and a population inversion was sustained, the brightness of the cloud can theoretically surge dramatically with the proper external stimulus.
In cases of powerful masers, brightness enhancements by factors of billions or more are feasible (see e.g. Elitzur 1992 and references therein). The observed increase in brightness is subject to various physical constraints, including the availability of excited atoms, the geometry of the cloud, and the duration of the pumping mechanism (see Figure 4).
The brightness of a cloud during a maser flare, relative to its equilibrium state, can increase dramatically because of the process of stimulated emission. In a maser, the presence of a population inversion allows radiation passing through the medium to be amplified exponentially. This occurs because photons stimulate the emission of additional photons, leading to a rapid multiplication of photon numbers and, consequently, a significant increase in intensity. The extent of this amplification is often described by an exponential factor, which is dependent on both the density of the inverted population and the path length through the cloud (Gray, 2012).
During the brief moment of population inversion, stimulated emission could dominate, leading to an amplified signal at the HI frequency. This could cause the cloud to exhibit a maser flare. The exact mechanism and theory behind this phenomenon will be investigated in a subsequent paper.
(a) Teegarden’s Star.
(b) Barnard’s Star.
(c) HD 157881.
(d) SGR 1935+2154.
(e) 3C 76.1 (B0300+162).
(f) 3C 365 (B1756+134).
(g) Void.
(h) Wow! Signal.
Figure 3: Hydrogen column density between -10 and +10 km/s near the observed targets (figures a to g) compared to the Wow! Signal region (figure h). The scan strip of our observations targets and the average density are shown as in the middle of the plot (white box and text). The Wow! Signal figure shows the beamsize of the two receivers used by the Big Ear. It is not clear in which of these receivers the signal was detected. Note that the colorbar is not on the same scale in each plot to highlight the relative column density. Data from the HI4PI Collaboration et al. (2016).

### v2

@@ 3 Arecibo SETI
Following the end of HRMS, Arecibo continued to participate in SETI through various initiatives, including the Search for Extraterrestrial Radio Emissions from Nearby Developed Intelligent Populations (SERENDIP) project (Werthimer et al., 1988). SERENDIP was a piggyback search, which means that it used data collected by Arecibo during other observations to perform a parallel search for extraterrestrial signals.
SERENDIP IV, which began in the late 1990s, used a multichannel receiver installed at Arecibo to scan the sky for signals over a wide range of frequencies (Chennamangalam et al., 2017). The data collected by SERENDIP IV were also used by the SETI@home project, a distributed computing effort that allowed millions of volunteers around the world to analyze SETI data on their home computers (Korpela et al., 2001). This collaboration between Arecibo and SETI@home significantly expanded the scope and reach of SETI research.
@@
Throughout its operational life, Arecibo was involved in numerous other SETI-related projects and observations, often serving as a testbed for new technologies and techniques in the search for extraterrestrial intelligence. The observatory’s unparalleled sensitivity and versatility made it a cornerstone of SETI research, and its contributions continue to influence the field even after its collapse in 2020.
The Arecibo Observatory’s legacy in SETI is marked by its role in pioneering early searches, conducting high-profile experiments like the Arecibo Message, and supporting a wide range of projects that advanced our understanding of the cosmos and the potential for life beyond Earth. The observatory data and findings remain valuable resources for ongoing and future SETI efforts (Tarter, 2001). The Big Ear and the Arecibo Observatory were two successful contemporary telescopes that also dedicated their time to SETI research (section 2).
@@ 4 Arecibo Wow!
In 2017 we started the Arecibo REDS project (Radio Emissions from Red Dwarf Stars). This project aims to study radio emissions from red dwarf stars using data from the Arecibo Observatory. Red dwarf stars, the most abundant type in the Milky Way, emit a wide range of radio frequencies because of their high activity levels (Yiu et al., 2024). We are studying these emissions, focusing on their frequency, polarization, and variability. Our observations are intended to help characterize stellar flares and coronal mass ejections, which impact the space weather and habitability of their planets (Airapetian et al., 2020).
Our research also explores the potential utility of these observations as tools for detecting exoplanets orbiting red dwarf stars (Vedantham et al., 2020). The results of Arecibo REDS will be presented elsewhere. After the collapse of the Arecibo Telescope in 2020, we plan to continue these observations with other radio telescopes.
@@
We observed from the Arecibo Observatory’s 305-meter telescope from 2017 to 2020 (AO Project A3123) and the 12-meter telescope in 2023 (AO Project A4006). Most of our targets were red dwarf stars with planets. Our observation protocol was similar to the one used by Route & Wolszczan (2012) to search for radio flares from ultracool dwarfs. Spectral observations were made in the L-band in full Stokes mode with up to 10-minute integration using the Arecibo Observatory’s Mock Spectrometer. Each of the seven receiver boxes had a bandwidth of 83 MHz at central frequencies of 1282, 1353, 1424, 1495, 1566, 1637, and 1708 MHz, for a total continuous bandwidth of 509 MHz.
Similar observations were made in the S, C, and X bands. The integration was 100 ms and the channel width 10 kHz, corresponding to a Doppler change of 2.1 km/s relative to the Hi line (unintentionally the same as the Big Ear during the Wow! Signal detection). There were 8192 channels per receiver. We used both diode and target calibrators to calibrate our flux density measurements.
Arecibo Wow! is a technosignatures side project of of Arecibo REDS. Our original observational strategy was neither tailored nor optimized for the detection of technosignatures, e.g. a high time and frequency resolution (Wright, 2021). However, our sample included stars with known potentially habitable exoplanets (PHL, 2024), which were observed for long sessions that spanned frequencies from 1 to 10 GHz.
In this study, we focus only on our drift-scan mode observations made from February to May 2020 at topocentric frequencies between 1419.5 and 1421.0 MHz (LABEL:tab:suppl-scans). This mode and frequency range are analogous to the protocol used in the detection of the Wow! Signal (section 2).
@@
The seven targets studied with drift scans were Teegarden’s Star, Barnard’s Star, and HD 157881, which were part of our main study (Table 4). We also conducted observations in this mode of the radio galaxies 3C 76.1 (B0300+162) and 3C 365 (B1756+134), which served as our calibration standards. The magnetar SGR 1935+2154 constituted an opportunistic target, due to its active state during the observation period and its proximal spatial alignment with our primary subjects (Bochenek et al., 2020). Our final target was a nearby random section of the sky that we denoted Void, for reference purposes.
@@ 5 Data Analysis and Results
The data were read and calibrated using standard procedures with the Arecibo Software Library. The dynamic spectra were detrended and normalized in frequency and time to cancel out broadband signals and identify narrowband variations. Most of the observations were made in the LSRK frame, but were also converted to the topocentric frame to make it easier to identify terrestrial RFI. Our longest drift scan was 10 minutes long (2.5 degrees in right ascension), of Teegarden’s star (Figure 2).
@@
Figure 2: Dynamic spectra of a 10-minute drift scan centered on Teegarden’s Star showing four narrowband signals associated to small cold Hi clouds, labeled A, B, C, and D. The top subplot is the time average flux density with the galactic Hi profile at the center. The right subplot is the frequency average flux density with a peak between 400 and 500 seconds attributable to AGC 74887. The clouds are not related to Teegarden’s Star or AGC 74887. Their properties are described in Table 5.
We developed a series of methodologies in the Interactive Data Language (IDL) aimed at the detection of narrowband signals. Among these, one emulated the analytical approach employed by the Ohio SETI, specifically averaging data over 10-second intervals and using the most recent 60 data points for mean and background noise calculation. The flux density was then converted to a signal-to-noise ratio (SNR). All of the methods yielded consistent results.
@@
We identified four main signals in the vicinity of Teegarden’s Star corresponding to small cold neutral hydrogen (Hi) clouds (Figure 2). Similar clouds appeared in the observations of SGR 1935+2154 and in the Void region. The intensity and location of the cloud signals were consistent between multiple independent scans during the same observation session of Teegarden’s Star. There was some variability in the SGR 1935+2154 signals observed in three epochs separated by a week, which could be attributed to noise or scintillation.
@@
As an example, we focus our analysis on the clouds near Teegarden’s Star, since it was the longest drift scan we performed. Their observed properties are consistent with small cold Hi clouds (Table 5 and LABEL:fig:suppl-clouds). Their hydrogen column density $n_{HI}$ corresponds to their partial contribution to the total Hi background (Figure 3). See the supplementary LABEL:fig:suppl-4HPI and LABEL:fig:suppl-GALFA for a description of $n_{HI}$ in all our observed targets.
(a) Teegarden’s Star (GALFA).
(b) Wow! Signal (GALFA).
[…]
The most noticeable characteristic of the Hi cloud signals depicted in Figure 2, is their close resemblance to the bandwidth and duration of the Wow! Signal (Figure 1). They are concentrated primarily in a bandwidth of $\approx$ 10 kHz, but it is clear that this is an artifact of the frequency resolution, and their actual bandwidth could be smaller or even slightly larger. These clouds are spatially resolved, among the characteristic galactic Hi profile, with different frequencies due to different Doppler velocities.
The angular dimensions of these Hi clouds are comparable to the azimuth beam width of the Big Ear telescope, approximately 8 arcminutes (section 2). Their brightness might fall within or below the detection sensitivity of the Big Ear. The primary distinction between these signals and the Wow! Signal is their brightness and transient characteristics, that is, they have to become about one-hundred times brighter for a few minutes to reproduce the Wow! Signal.